# Improved Counting under Continual Observation with Pure Differential Privacy

### v1

@@ 3.2 $k$-ary Tree Mechanisms
Additionally, let $\Sigma(I)=\sum_{i\in I}\mathbf{x}_{i}$ for $I\in\mathcal{T}_{k}$ when such a tree is used to store partial sums on $\mathbf{x}$, $\widehat{\Sigma}$ for the case where a vertex in the tree stores a prefix sum plus noise. Also, we will find it expedient to express integers using $k$-ary digits, which we define next.
@@
As Theorem 3.7 is not our main contribution, we provide a sketch of proof in A.4. Before continuing and including subtraction as a technique, we note that Theorem 3.7 describes the standard binary tree mechanism for $k=2$, and that the leading constant in the error is minimized for $k=17$, in which case it equals $\approx 0.234$.
@@ 3.3 $k$-ary Trees With Subtraction
Now, we only consider the case of odd $k$ for a few reasons: (1) if $k$ is odd, then we have symmetry in the digits, meaning we add at worst as many things as we may subtract. This makes the analysis easier. (2) it will turn out to be the case that even $k$ is not optimal, see Section A.3 for more details. (3) if $k$ is odd, then there is a simple relation for when we use the root vertex of a tree or not to produce an output: we only use the root vertex if $t$ is in the right half of the tree. This gives an intuitive condition on the height to support $T$ inputs: $h=\lceil\log_{k}(2T)\rceil$ – set the height such that there is a sufficient number of left-half leaves. This can be compared to standard $k$-ary trees, where the same condition is to use every leaf of the tree except the right-most one.
@@ 4.3 Space and time complexity
5 $\varepsilon$-DP vs. $(\varepsilon,\delta)$-DP mechanisms for continual counting
Any $\varepsilon$-DP mechanism can achieve a smaller mean squared error, up to lower order asymptotically vanishing terms, than all known Gaussian noise $(\varepsilon,\delta)$-DP mechanisms, assuming $\delta$ is picked small enough. Lemma 5.1 formalizes this for the case of such mechanisms with the best known error scaling.
@@
For constants $C_{\varepsilon},C_{\varepsilon,\delta}>0$, let $(C_{\varepsilon}/\varepsilon^{2})\log(T)^{3}+o\big{(}\log(T)^{3}\big{)}$ be a bound on the mean squared error of an $\varepsilon$-DP mechanism, and $(C_{\varepsilon,\delta}/\varepsilon^{2})\log(T)^{2}\log(1/\delta)+o\big{(}\log(T)^{2}\log(1/\delta)\big{)}$ be a bound on the mean squared error of an $(\varepsilon,\delta)$-DP mechanism, both for continual counting. Then the $\varepsilon$-DP mechanism achieves a smaller bound, up to lower-order asymptotically vanishing terms, on the mean squared error for $\delta=O(1/T^{\alpha})$ where $\alpha=C_{\varepsilon}/C_{\varepsilon,\delta}$.
@@
We want to emphasize that this is in and of itself not surprising – rather, what may be surprising, is what this assumption on $\delta$ translates to for actual constants $C_{\varepsilon}$ and $C_{\varepsilon,\delta}$. To this end, we re-state a result from [19] (Theorem 1.1 and 1.2) below as Theorem 5.3.
Theorem 5.3 (Optimal approximate-DP factorization mechanism [19]).
Refer to the class of $(\varepsilon,\delta)$-DP mechanism based on factorization and using Gaussian noise as $\mathbb{M}$. There is a mechanism in $\mathbb{M}$ that achieves a bound on the mean squared error of $\frac{4}{\varepsilon^{2}\pi^{2}\log(e)^{3}}\log(T)^{2}\log(1/\delta)+o\big{(}\log(T)^{2}\log(1/\delta)\big{)}$, and any further improvement in the leading constant is only possible through a tighter analysis of the Gaussian mechanism.
This is quite a strong statement as all currently known, unbiased, mechanisms for continual counting in $(\varepsilon,\delta)$-DP, with $\delta>0$, belongs to $\mathbb{M}$ [19]. Identifying that Algorithm 2 has a leading constant of $C_{\varepsilon}=0.124$ for $k=19$, and that the corresponding constant in Theorem 5.3 is $C_{\varepsilon,\delta}=4/(\pi^{2}\log(e)^{3})=0.135$, using Lemma 5.1 we arrive at $\delta=O(T^{-0.92})$. This holds for other $\varepsilon$-DP mechanisms too, but with stricter conditions on $\delta$. The only caveat here, also stated in Theorem 5.3, is that the constant given in [19] is valid for all $10<\varepsilon,\delta<$ 1. It seems likely that the constant be reduced via a tighter analysis of the Gaussian mechanism, as done in for example [2].
@@ 6 Discussion
While there are no fundamental reasons ruling out asymptotic improvements in the mean squared error for $\varepsilon$-DP ([19] provides an $\Omega(\log(T)^{2})$ lower bound), there have been no asymptotical improvements since the binary tree mechanism [13, 5]. This is the context of our paper, improving constants, and we make no claim that no further constant improvements are possible. We do make a note of, however, that Algorithm 2 is easy to implement, efficient and offers the best known error-scaling to date.
To further improve on the error, one could therefore imagine that using the techniques in [1] could reduce constants slightly, by only using a subset of the leaves to lower the sensitivity. However, the fraction $1/k$ of unused vertices is small for our order of $k$, meaning the improvement in sensitivity would be small, and if we are using a subset of leaves, then the structure of the queries that we need to average over are hard to analyze, unlike in the case for $k=2$. So this technique might reduce constants further, but not by much, and it would be non-trivial to show. A more promising direction would likely involve searching among factorization mechanisms directly, as done in [19, 16, 11], but instead for $\varepsilon$-DP.
Appendix A Appendix
A.1 Definitions for Differential Privacy
@@
A.2 $k$-ary Tree Mechanisms: Proofs
@@
Proof A.4 (Proof sketch of Theorem 3.7).
@@
A.3 Leveraging Subtraction For Even $k$
@@
First note that for even $k$ we lose the symmetry in the offset digits, so we need to make a decision if the digits in the offset representation are $\{-k/2+1,\dots k/2\}$ or $\{-k/2,\dots,k/2-1\}$. We pick the first one as that allows for supporting more inputs, giving Definition A.5.
Definition A.5 (Offset (even) $k$-ary representation of integers).
@@
Substituting the representation used in Algorithm 2 by the one in Definition A.5 and setting the height appropriately gives a valid algorithm using a $k$-ary tree. However, it performs worse than using odd $k$.
@@
Proof A.6 (Proof sketch.).

### v2

@@ 3.2 $k$-ary Tree Mechanisms
Additionally, let $\Sigma(I)=\sum_{i\in I}\mathbf{x}_{i}$ for $I\in\mathcal{T}_{k}$ when such a tree is used to store partial sums on $\mathbf{x}$, $\widehat{\Sigma}$ for the case where a vertex in the tree stores a partial sum plus noise. Also, we will find it expedient to express integers using $k$-ary digits, which we define next.
@@
As Theorem 3.7 is not our main contribution, we provide a proof sketch in B.1. Before continuing and including subtraction as a technique, we note that Theorem 3.7 describes the standard binary tree mechanism for $k=2$, and that the leading constant in the error is minimized for $k=17$, in which case it equals $\approx 0.234$.
@@ 3.3 $k$-ary trees with subtraction
Now, we only consider the case of odd $k$ for a few reasons: (1) if $k$ is odd, then we have symmetry in the digits, meaning we add at worst as many things as we may subtract. This makes the analysis easier. (2) it will turn out to be the case that even $k$ is not optimal, see Appendix C for more details. (3) if $k$ is odd, then there is a simple relation for when we use the root vertex of a tree or not to produce an output: we only use the root vertex if $t$ is in the right half of the tree. This gives an intuitive condition on the height to support $T$ inputs: $h=\lceil\log_{k}(2T)\rceil$ – set the height such that there is a sufficient number of left-half leaves. This can be compared to standard $k$-ary trees, where the same condition is to use every leaf of the tree except the right-most one.
@@ 4.3 Space and time complexity
5 Laplace vs Gaussian noise
Having designed an efficient $\varepsilon$-DP algorithm for continual counting with a state-of-the-art bound on the error, a natural question arises: how does it compare to $(\varepsilon,\delta)$-DP mechanisms?
Since the noise of the Gaussian mechanism grows as $\delta$ approaches zero, it is clear that for any given bounds on $\ell_{1}$ and $\ell_{2}$ sensitivity, using the Gaussian mechanism with sufficiently small $\delta$ will result in more noise than the Laplace mechanism. We will see that for current continual counting methods this change happens for surprisingly large values of $\delta$, with the exact threshold depending on the number of time steps $T$.
Furthermore, we adapt the Laplace mechanism to work with $\ell_{2}$-sensitivity and show that it provides a meaningful $(\varepsilon,\delta)$-DP guarantee with variance close to that of the Gaussian mechanism. For proofs not included we refer the reader to Appendix D.
5.1 Laplace vs Gaussian noise for continual counting
We start by stating a lemma that compares the currently best continual counting mechanisms under pure and approximate differential privacy.
@@ 5.1 Laplace vs Gaussian noise for continual counting
For constants $B_{\varepsilon},B_{\varepsilon,\delta}>0$, let $(B_{\varepsilon}/\varepsilon^{2})\log(T)^{3}+o\big{(}\log(T)^{3}\big{)}$ be a bound on the mean squared error of an $\varepsilon$-DP mechanism, and $(B_{\varepsilon,\delta}/\varepsilon^{2})\log(T)^{2}\log(1/\delta)+o\big{(}\log(T)^{2}\log(1/\delta)\big{)}$ be a bound on the mean squared error of an $(\varepsilon,\delta)$-DP mechanism, both for continual counting. Then the $\varepsilon$-DP mechanism achieves equivalent or better mean squared error, up to lower-order asymptotically vanishing terms, for $\delta=O(1/T^{B_{\varepsilon}/B_{\varepsilon,\delta}})$.
@@
Proof 5.3 (Proof of Theorem 1.2).
Algorithm 2 has a leading constant of $B_{\varepsilon}<0.124$ for $k=19$ under $\varepsilon$-DP, and the corresponding constant in [24] is $B_{\varepsilon,\delta}=4/(\pi^{2}\log(e)^{3})>0.1349$, so we have $B_{\varepsilon}/B_{\varepsilon,\delta}<0.92$. Lemma 5.1 thus implies that the pure mechanism has a smaller bound on the noise for $\delta=O(T^{-0.92})$.
5.2 Laplace noise scaled to $\ell_{2}$-sensitivity
We next consider how we may extend the Laplace mechanism with an approximate DP guarantee similar to the Gaussian mechanism, i.e., with noise scaled to the $\ell_{2}$-sensitivity. Since the proof of this is straightforward using known results we suspect this fact may be known, but we are not aware of it appearing in the literature.
Lemma 5.4 ($\ell_{2}$ Laplace Mechanism).
Let $f:\mathcal{X}^{n}\to\mathbb{R}^{d}$ be a function with $\ell_{2}$-sensitivity $\Delta_{2}\coloneqq\max_{\mathcal{D}\sim\mathcal{D}^{\prime}}\left\lVert f(\mathcal{D})-f(\mathcal{D}^{\prime})\right\rVert_{2}$. For a given dataset $\mathcal{D}\in\mathcal{X}^{n}$ and $\varepsilon,\delta\in(0,1)$, the output $f(\mathcal{D})+\mathsf{Lap}(\Delta_{2}/a_{\varepsilon,\delta})^{d}$, where $a_{\varepsilon,\delta}=\sqrt{2\ln(1/\delta)}(\sqrt{1+\varepsilon/\ln(1/\delta)}-1)$, satisfies $(\varepsilon,\delta)$-DP.
The variance needed for using Laplace noise instead of Gaussian noise for $(\varepsilon,\delta)$-DP is only a constant factor higher whenever $\varepsilon\leq\log(1/\delta)$, and the constant approaches $2$ as $\varepsilon$ approaches zero.
Corollary 5.5.
Let $\mathbf{y}^{\mathrm{lap}}$ and $\mathbf{y}^{\mathrm{gauss}}$ be outputs based on the $\ell_{2}$ Laplace mechanism (Lemma 5.4) and the Gaussian mechanism (Lemma A.3) respectively, both to achieve the same $(\varepsilon,\delta)$-DP guarantee where $\varepsilon\leq\ln(1/\delta)$. Then $\operatorname{Var}[\mathbf{y}^{\mathrm{lap}}_{j}]\leq\frac{(\varepsilon/\ln(1/\delta))^{2}}{2(\sqrt{1+\varepsilon/\ln(1/\delta)}-1)^{2}}\cdot\operatorname{Var}[\mathbf{y}^{\mathrm{gauss}}_{j}]$. Also, $\lim_{\varepsilon\to 0}\frac{(\varepsilon/\ln(1/\delta))^{2}}{2(\sqrt{1+\varepsilon/\ln(1/\delta)}-1)^{2}}=2$.
Note that while functions with low $\ell_{1}$-sensitivity also have low $\ell_{2}$-sensitivity, the converse does not hold. In principle we can replace Gaussian noise with Laplace noise at a limited loss in utility for approximate DP (Corollary 5.5), but the corresponding $\varepsilon$-DP guarantee afforded by the noise will be weak unless $\Delta_{1}$ is small. Theorem 1.3 can alternatively be seen as adding an approximate DP guarantee to pure DP mechanisms built on top of the Laplace mechanism.
@@ 6 Discussion
While there are no known fundamental reasons ruling out asymptotic improvements in the mean squared error for $\varepsilon$-DP, it might be possible to match the $\Omega(\log(T)^{2})$ lower bound of [24], there have been no asymptotic improvements since the binary tree mechanism [16, 5]. This state of affairs motivates exploring how much leading constants can be improved using the factorization mechanism framework. Algorithm 2 falls into the framework, is easy to implement, efficient, and offers the best known error-scaling to date.
Our paper also reveals that using Laplace noise can improve known “optimal” factorization mechanisms using Gaussian noise under rather weak assumptions on $\delta$. An interesting future direction is to investigate, for fixed $\varepsilon,\delta$ and a given factorization: which noise distribution achieves optimal error? It seems unlikely that Laplace or Gaussian noise is always best – indeed, Vinterbo [37] has results of this flavor in more restricted settings.
Appendix A Definitions for Differential Privacy
@@
Appendix B $k$-ary Tree Mechanisms: Proofs
@@
Proof B.1 (Proof sketch of Theorem 3.7).
@@
Appendix C Leveraging subtraction for even $k$
@@
First note that for even $k$ we lose the symmetry in the offset digits, so we need to make a decision if the digits in the offset representation are $\{-k/2+1,\dots k/2\}$ or $\{-k/2,\dots,k/2-1\}$. We pick the first one as that allows for supporting more inputs, giving Definition C.1.
Definition C.1 (Offset (even) $k$-ary representation of integers).
@@
Substituting the representation used in Algorithm 2 by the one in Definition C.1 and setting the height appropriately gives a valid algorithm using a $k$-ary tree. However, it performs worse than using odd $k$.
@@
Proof C.2 (Proof sketch.).
@@
Appendix D Laplace vs Gaussian noise: Proofs
Proofs missing from the Section 5 can be found here. The technique for extending the Laplace mechanism to approximate DP relies on the following heterogeneous composition theorem [29].
Lemma D.1 (Simplified version of Theorem 3.5 in [29]).
For any $\varepsilon_{j}>0$ for $j\in\{1,,\dots,k\}$, and $\delta\in(0,1)$, the class of $\varepsilon_{j}$-differentially private mechanisms satisfy $(\tilde{\varepsilon},\delta)$-differential privacy under $k$-fold adaptive composition for
$\tilde{\varepsilon}=\min\bigg{\{}\sum_{j=1}^{k}\varepsilon_{j},\,\sum_{j=1}^{k}\frac{(e^{\varepsilon_{j}}-1)\varepsilon_{j}}{e^{\varepsilon_{j}}+1}+\sqrt{2\ln(1/\delta)\sum_{j=1}^{k}\varepsilon_{j}^{2}},\bigg{\}}\enspace.$
Proof D.2 (Proof of Lemma 5.4).
For $x_{1},\dots,x_{j}\in(0,1)$, since $\frac{e^{x_{i}}-1}{e^{x_{i}}+1}\leq x_{i}/2$ it follows from Lemma D.1 that $(\tilde{\varepsilon},\delta)$-differential privacy is also satisfied for
$\tilde{\varepsilon}=\min\bigg{\{}\sum_{j=1}^{k}\varepsilon_{j},\,\frac{1}{2}\sum_{j=1}^{k}\varepsilon_{j}^{2}+\sqrt{2\ln(1/\delta)\sum_{j=1}^{k}\varepsilon_{j}^{2}},\bigg{\}}\enspace.$ (1)
It suffices to establish the privacy of the Laplace mechanism when restricted to arbitrary neighboring inputs $x$, $x^{\prime}$. We will reason about the privacy loss for each coordinate of $\mathbf{y}=f(x)$ separately relative to the neighboring output $\mathbf{y}^{\prime}=f(x^{\prime})$. Focusing on one coordinate $\mathbf{y}_{j}$ and its private release, we have that its release must satisfy $\varepsilon_{j}$-differential privacy where $\varepsilon_{j}=(a_{\varepsilon,\delta}/\Delta_{2})\cdot\lvert\mathbf{y}_{j}-\mathbf{y}_{j}^{\prime}\rvert$. Composing the $d$ private releases of each coordinate using the bound (1) and noting that $\sum_{j=1}^{d}\varepsilon_{j}^{2}\leq(a_{\varepsilon,\delta}/\Delta_{2})^{2}\sum_{j=1}^{d}(\mathbf{y}_{j}-\mathbf{y}_{j}^{\prime})^{2}\leq a_{\varepsilon,\delta}^{2}$ therefore gives an $(\varepsilon,\delta)$-guarantee where $\varepsilon=a_{\varepsilon,\delta}(\frac{a_{\varepsilon,\delta}}{2}+\sqrt{2\ln(1/\delta)})$. Proceeding to solve for (positive) $a_{\varepsilon,\delta}$ gives the statement.
Proof D.3 (Proof of Corollary 5.5).
Using Lemma 5.4 and noting that the variance of the noise added is twice the square of the scale parameter for Laplace noise, we get that
$\frac{\operatorname{Var}[\mathbf{y}^{\mathrm{lap}}_{j}]}{\operatorname{Var}[\mathbf{y}^{\mathrm{gauss}}_{j}]}=\frac{2\Delta_{2}^{2}/(2\ln(1/\delta)(\sqrt{1+\varepsilon/\ln(1/\delta)}-1)^{2})}{2\ln(1.25/\delta)\Delta^{2}_{2}/\varepsilon^{2}}\leq\frac{(\varepsilon/\ln(1/\delta))^{2}}{2(\sqrt{1+\varepsilon/\ln(1/\delta)}-1)^{2}}\enspace.$
To show the limit, note that for positive $w\leq 1$ we have that $\frac{w}{\sqrt{1+w}-1}\leq\frac{w}{w/2-w^{2}/8}=\frac{2}{1-w/4}\to 2$ as $w\to 0$, and $\frac{w}{\sqrt{1+w}-1}\geq 2$, both based on truncated Taylor expansions of $\sqrt{1+w}$.
Proof D.4 (Proof of Theorem 1.3).
From the proof of Lemma 5.4, we have that adding noise $\mathsf{Lap}(\Delta_{2}/a)$ gives an $(\varepsilon^{\prime},\delta)$-guarantee where $\varepsilon^{\prime}=a(\frac{a}{2}+\sqrt{2\ln(1/\delta)})$ if $\varepsilon^{\prime},\delta\in(0,1)$. Note that our setup is the same if we set $a=\frac{\Delta_{2}}{\lambda}$, and that adding noise $\mathsf{Lap}(\lambda)$ gives an $\frac{\Delta_{1}}{\lambda}$-DP guarantee. We can therefore always guarantee $(\varepsilon,\delta)$-DP where $\varepsilon=\min\{\frac{\Delta_{1}}{\lambda},\varepsilon^{\prime}\}<1$ if we enforce $\lambda>\Delta_{1}$.